P6-10 LO 6.6 Factoring Accounts Receivable Faeber Textile Company frequently factors its accounts receivable. During 2019, Fa| || |
Answer #1

Answer :

From the given information the journal entries should be calculated as :

Particulars Calculation Amount
Debit Credit
To record sale of goods on credit :
A/c receivable $ 100,000
Sales $ 100,000
To record collection of factored of accounts receivable :
Cash

= [70,000 * 90%] - 8,400

= [70,000 * 0.90] - 8,400

= 63,000 - 8,400

= 54,600

$ 54,600

Factoring expenses

= 70,000 * 12%

= 70,000 * 0.12

= 8,400

$ 8,400
Receivable from factor

= 70,000 * 10%

= 70,000 * 0.10

= 7,000

$ 7,000
A/c receivable

= 54,600 + 8,400 + 7,000

= 70,000

$ 70,000
To record collection of un-factored accounts receivable :
Cash

= 30,000 - 576

= 29,424

$ 29,424
Sales discount

= 30,000 * 96% * 2%

= 30,000 * 0.96 * 0.02

= 576

$ 576
A/c receivable

= 100,000 - 70,000

= 30,000

$ 30,000
To record collection of balance factored receivable :
Cash

= 7,000 - 132

= 6,868

$ 6,868
Sales discount

= 7000 * 94% * 2%

= 132

$ 132
Receivable from factor

= 70,000 - [90%*70,000]

= 7,000

$ 7,000